​Aki's Bicycle Designs has determined that when x hundred bicycles are​ built, the average cost per bicycle is given by​ C(x)=0.5x^2-1.4x-1+2.134​, where​ C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per​ bicycle?
Answer by mananth(16946)

C(x)=0.5x^2-1.4x-1+2.134​,
find the vertex
this is true when
b=-1.4, a=0.5
x = -b/(2a)
x = -(-1.4)/(2*0.5)
x = (1.4)/(1)
x = 1.4
.
solution: 1.4*100 =140 bicycles